# Analysis of Emotion in Rumour Threads on Social Media

### v1

@@ Introduction
Then we visualize it using a heatmap, e.g. in Figure 5, the cell 0.55 in the last row of the third column has is dark red in color, indicating that the emotion transition pair (love, fear) appears more frequently in rumour than non-rumour comments in Twitter15. Overall, we observe larger emotion transition probability mass in positive–positive and negative–negative emotion transitions. This indicates that emotions are contagious, aligning with psychological findings (Goldenberg and Gross 2019; Herrando and Constantinides 2021). Contrasting rumour and non-rumour comments, we observe common patterns, namely that fear–fear and love–sadness are more common in rumour comments, and love–joy and love–optimism appear more frequently in non-rumour comments. We also see differences among datasets: Twitter15 has more anger response to almost all emotions more in non-rumour posts; Twitter16 has a lot of anger and disgust in response to positive emotions in rumours. We aggregate emotions into Negative, Neutral and Positive emotions in Table 4. We observe positive values in emotion pairs where the transition ends with a negative emotion, indicating that discussions in rumours often trigger negative responses. On the contrary, negative delta values are observed in PHEME, Twitter16, suggesting non-rumours tend to prompt more positive responses.
@@
Emotion Trajectory
Figures 7, 8 and 9 illustrate the cumulative emotion across each dataset over time. At each chronological step, the counts represent the total number of observed emotions. Generally, we see a strong linear trend across datasets for all emotions. To better capture the rate of growth for each emotion, we apply linear regression and present the slopes in Table 5. From the table, it is apparent that negative emotions tend to grow faster in rumour posts than in non-rumour posts across all datasets, while positive emotions grow faster in non-rumour posts.
Causal Analysis
The causal relationships revealed in Figure 10 demonstrate several key patterns. First, we find out that the fact that a given thread is a rumour is not directly connected with other emotions. Specifically, the rumour has to rely on the emotion of surprise as a bridge to interact with other emotions, namely, $rumour\not\perp\!\!\!\perp Fear|Surprise$ and $rumour\not\perp\!\!\!\perp Anticipation|Surprise$. The change in the distribution of the rumour does not influence other emotions except surprise. This aligns with cognitive basis of rumour transmission, where surprising or counterintuitive information tends to capture attention and facilitate rumour spreading (Knapp 1944; Āllport and Postman 1947). Second, pessimism is primarily influenced by negative emotions (sadness and fear), while optimism is causally influenced by positive emotions (joy, love, and trust). Notably, there’s undirected edge between anger and disgust, this relationship aligns with our previous findings that both rumour and non-rumour posts exhibit intense expressions of these emotions.
@@
There are also a few counterintuitive findings, including sadness leading to joy, joy causing pessimism, and the causal relationship between disgust and love. We conducted a qualitative analysis of the 50 samples and found there are several possible reasons for this: (1) there is a complex interplay of emotions in social media interactions, where emotional responses are shaped by context and individual perspectives. For example, we had one response where joy was detected, “this tweet gives me hope that she may write an eighth” to the post “once again, jk rowling is not working on an eighth harry potter book.” where sadness is detected. Sadness can sometimes lead to joy when people use humor or shared memories to find solace in sadness, which serves as a coping mechanism for processing uncomfortable or shocking topics. Similarly, expressions of joy can paradoxically evoke pessimism in certain contexts, as the same post can be interpreted in vastly different ways depending on the readers’ emotional state, cultural background, or personal experiences. The dual causal relationship between disgust and love further emphasizes the complexity of emotions expressed in text. A post that initially provokes disgust might also elicit admiration or affection when audiences recognize an underlying message of authenticity, vulnerability, or humor. (2) Moreover, the analysis reveals challenges in accurately interpreting emotions through automatic labeling methods. Sarcasm and humor are frequently misclassified, with sarcasm often mistaken for joy due to its seemingly optimistic wording. The lack of contextual information leads to noise and inaccuracies in emotional categorizations. (3) The complexity of social media participants also contributes to this. Some users engage with posts for self-serving purposes, such as promoting their brand or gaining visibility, rather than genuinely responding to the content. These interactions together add a layer of noise to the results, making it even more challenging.
Related Work
The definition of rumour is generally complicated and varies from one publication to another. Some early work treated rumour as information that is false (Cai, Wu, and Lv 2014). Most recent definitions of rumours are “unverified and instrumentally relevant information statements in circulation” (DiFonzo and Bordia 2007) and “unverified information at the time of the posting”. This definition also aligns with the concept in recent work (Zubiaga et al. 2018, 2015; Tian, Zhang, and Lau 2022) and the Oxford English Dictionary, which defines the rumour as “an unverified or unconfirmed statement or report circulating in a community”. [^2] https://www.oed.com/dictionary/rumour_n?tab=meaning_and_use
Existing research highlights the significant role of emotions in understanding general misinformation, mostly fake news. Research has found relationships exist between negative sentiment and fake news, and between positive sentiment and genuine news (Zaeem et al. 2020). Fake news also expresses a higher level of overall emotion, negative emotion, and anger than real news (Zhou, Tao, and Zhang 2022). Negative emotions like sadness and anger can serve as indicators of misinformation (Prabhala and Bose 2019). The role of emotions in rumours has been recognized since the Second World War, reflecting the interactive and community-driven nature of rumour spreading. Knapp’s taxonomy (Knapp 1944) of rumours categorizes them into three types, each deeply embedded with emotions: (1) ‘pipedream’ rumours, which evoke wishful thinking; (2) ‘bogy’ rumours, which heighten anxiety or fear; and (3) ‘wedge-driving’ rumours, which incite hatred. This taxonomy underscores how rumours are inherently embedded with emotional undercurrents.”
Recent research on emotion in rumours largely focuses on their role in spreading behaviour, some studies have used questionnaires to gather participants’ reactions to specific rumours (Zhang et al. 2022; Rijo and Waldzus 2023; Ali et al. 2022), while others have employed cascade size and lifespan as indicators (Pröllochs, Bär, and Feuerriegel 2021b, a). Key findings of such work include: rumours conveying anticipation, anger, trust, or offensiveness tend to generate more shares, have longer lifespans, and exhibit higher virality (Pröllochs, Bär, and Feuerriegel 2021b). Additionally, false rumours containing a high proportion of terms reflecting positive sentiment, trust, anticipation, anger, or condemnation are more likely to go viral (Solovev and Pröllochs 2022; Pröllochs, Bär, and Feuerriegel 2021a). However, existing research has notable gaps: it often focuses on isolated aspects of emotions in rumours, primarily identifies correlations rather than causality, and tends to examine rumour data in isolation. To address these gaps, we aim to propose a comprehensive emotional analytical framework that integrates multiple emotion-related tasks, contrasting rumour and non-rumour content with the aim to enhance our understanding of emotions and, ultimately, to improve rumour detection in online social media.
@@ Conclusion
In this work, we presented a analytical emotion framework for online rumours. We make the use of EmoLLM for automatic emotion detection and ground its predictions to the categorical emotion labeling scheme. The framework analyzes the emotion from various aspects, from simply emotion polarity (sentiment valence), emotion distribution, emotion transition and trajectory, we also analyze the causal relationship between rumour label and emotions. The key findings include: compared with non-rumour contents, rumour are significantly more negative in sentiments, containing more negative emotions like anger, fear and pessimism; emotions are contagious in online social context, rumour contents usually trigger negative responses and non-rumours tend to receive positive ones; Cumulative emotion regression coefficient showed that negative emotions grow significantly faster in rumours comments as positive emotions in non-rumour ones; the rumour tweets are not directly connected with other emotions and rely on the emotion surprise as a bridge. Pessimism is primarily influenced by negative emotions (sadness and fear), while optimism is causally influenced by positive emotions (joy, love, and trust), anger and disgust exhibit bidirectional causation. By presenting the framework, we hope to facilitate research in more comprehensive and fine-grained study in emotion in online rumour contents and better detection techniques.
@@
This work also faces several challenges and limitations: 1) We rely on EmoLLM as our automatic emotion detection tool for all emotion-related tasks. While it is generally efficient and effective, it exhibits inaccuracies in analyzing complex online discussions, such as those involving sarcasm; 2) Although we have access to the chronological order of tweets within conversations, explicit conversation structures are not available for all data; and 3) The datasets used in this study are limited to English textual rumour data. Future work should explore multilingual and multimodal content in rumour conversations to provide a more comprehensive analysis.
@@
is under CC-BY license, Twitter15 and Twitter 16444https://github.com/majingCUHK/rumour_RvNN are under MIT License, EmoLLM [^5] https://github.com/lzw108/EmoLLMs is under MIT License and Huggingface Transformers [^6] https://github.com/huggingface/transformers is under Apache License 2.0). Our source code and annotated data will be under MIT license.

### v2

@@ Introduction
Then we visualize it using a heatmap, e.g. in Figure 5, the cell with a value of 0.55 in the last row of the third column is dark red, indicating that the emotion transition pair (love, fear) appears more frequently in rumour than non-rumour comments in Twitter15. Overall, we observe larger emotion transition probability mass in positive–positive and negative–negative emotion transitions.
This indicates that emotions are contagious, aligning with psychological findings (Goldenberg and Gross 2019; Herrando and Constantinides 2021). Contrasting rumour and non-rumour comments, we observe common patterns, namely that fear–fear and love–sadness are more common in rumour comments, and love–joy and love–optimism appear more frequently in non-rumour comments. We also see differences among datasets: Twitter15 has more anger response to almost all emotions more in non-rumour posts; Twitter16 has a lot of anger and disgust in response to positive emotions in rumours. We aggregate emotions into Negative, Neutral and Positive emotions in Table 4. We observe positive values in emotion pairs where the transition ends with a negative emotion, indicating that discussions in rumours often trigger negative responses. On the contrary, negative delta values are observed in PHEME, Twitter16, suggesting non-rumours tend to prompt more positive responses.
Emotion Trajectory
Figures 7, 8 and 9 illustrate the cumulative emotion over time in PHEME, Twitter15 and Twitter 16 results. At each chronological step, the counts represent the total number of observed emotions. Generally, we see a strong linear trend across datasets for all emotions. To better capture the rate of growth for each emotion, we apply linear regression and present the slopes in Table 5. From the table, it is apparent that negative emotions tend to grow faster in rumour posts than in non-rumour posts across all datasets, while positive emotions grow faster in non-rumour posts.
@@
Algorithm 1 Emotion Causal Relationship Discovery
1:Input: Rumour Threads $\mathbf{X}$, significance level $\alpha$
2:Output: Completed Partially Directed Acyclic Graph (CPDAG)
4:Step 1: Skeleton Identification
5:for each pair of variables $(X,Y)$ in $G$ do
7:     if such a subset $S$ exists then
8:         Remove the edge $X-Y$ from $G$.
9:     end if
10:end for
11:Step 2: Edge Orientation
13:     if $Z\notin S$ for all separating sets $S$ for $X$ and $Y$ then
14:         Orient as $X\to Z\leftarrow Y$ (identify a collider).
15:     end if
16:end for
17:while possible do
18:     for each edge $(X-Y)$ in $G$ do
19:         if there exists a directed path $X\to\dots\to Z$ such that $Z-Y$ then
20:              Orient as $X\to Y$ (acyclicity rule).
21:         else if orienting $X-Y$ as $X\to Y$ creates a new v-structure then
22:              Orient as $X\to Y$ (v-structure rule).
23:         end if
24:     end for
25:end while
Causal Relationship of Emotions
To gain a deeper insight into the relationship between rumours and the emotions underlying them, we extend our analysis beyond statistical correlation by conducting a causal analysis. Specifically, we apply the Peter-Clark (PC) algorithm (Spirtes, Glymour, and Scheines 2000), a classical constraint-based causal discovery algorithm on the merged of PHEME, Twitter15 and Twitter16 datasets.
Under the fundamental assumption of causal Markov condition that a variable is conditionally independent of all its non-effects given its direct cause, faithfulness ensures that the casual graph exactly encodes the independence and conditional independence relations among emotion and rumour label variables. These two assumptions allow us to infer causal relationships from observed statistical independencies, forming the cornerstone of constraint-based causal discovery methods. The PC algorithm identifies causal relationships among the variables of interest, represented as a directed acyclic graph (DAG), by numerating the independence and conditional independence relationships. The algorithm consists of two main steps:
  1. 1.
  2. 2.
    Edge Orientation: After constructing the skeleton, edges are oriented by a set of predefined rules, Meek’s Rule (Meek 1997) to avoid cycles and orient collider structures.
The complete PC algorithm is provided in Algorithm 1. It returns a completed partially directed acyclic graph of emotions in the thread, which represents an equivalence class of causal graphs that are consistent with the observed data independence and conditional independence relations. In our implementation, we adopt the Fisher-z test (Fisher 1921) to infer the conditional independence relations.
The causal relationships revealed in Figure 10 demonstrate several key patterns. First, we find out the fact that a given thread is a rumour is not directly connected with other emotions. Specifically, the rumour has to rely on the emotion of surprise as a bridge to interact with other emotions, namely, $rumour\not\perp\!\!\!\perp Fear|Surprise$ and $rumour\not\perp\!\!\!\perp Anticipation|Surprise$. The change in the distribution of the rumour does not influence other emotions except surprise. This finding aligns with cognitive basis of rumour transmission in previous work, where surprising, novel or counterintuitive information tends to capture attention and facilitate rumour spreading (Knapp 1944; Irving 1947; Vosoughi, Roy, and Aral 2018). Second, pessimism is primarily influenced mostly by negative emotions (sadness and fear), while optimism is causally influenced by positive emotions (joy, love, and trust). Notably, there’s undirected edge between anger and disgust, this relationship aligns with our previous findings that both rumour and non-rumour posts exhibit intense expressions of these emotions.
@@
There are also a few counterintuitive findings, including sadness leading to joy, joy causing pessimism, and the causal relationship between disgust and love. We conducted a qualitative analysis of the 50 samples and found there are several possible reasons for this: (1) There is complex interplay of emotions in social media interactions, where emotional responses are shaped by context and individual perspectives. For example, we had one response where joy was detected, “this tweet gives me hope that she may write an eighth” to the post “once again, jk rowling is not working on an eighth harry potter book.” where sadness or pessimism is detected. Sadness can sometimes lead to joy when people use humor or shared memories to find solace in sadness, which serves as a coping mechanism for processing uncomfortable or shocking topics. Similarly, expressions of joy can paradoxically evoke pessimism in certain contexts, as the same post can be interpreted in vastly different ways depending on the readers’ emotional state, cultural background, or personal experiences. The undirected edge between disgust and love further emphasizes the complexity of emotions expressed in text. A post that initially provokes disgust might also elicit admiration or affection when audiences recognize an underlying message of authenticity, vulnerability, or humor. (2) There are prediction errors. Sarcasm and humor are frequently misclassified, with sarcasm often mistaken for joy due to its seemingly optimistic wording. The lack of contextual information leads to noise and inaccuracies in emotional categorizations. (3) Social media interaction can be unpredictable. Some users engage with posts for self-serving purposes, such as promoting their brand or gaining visibility, rather than genuinely responding to the content.
@@ Conclusion
In this work, we presented a analytical emotion framework for online rumours. We make the use of EmoLLM for multi-aspects affective information annotation and analysis. The framework analyzes the emotion from direct emotion polarity (sentiment valence), emotion distribution, emotion transition and trajectory to causal relationship between rumour and emotions. The key findings include: compared with non-rumour contents, rumour are significantly more negative in sentiments, containing more negative emotions like anger, fear and pessimism; emotions are contagious in online social context, rumour contents usually trigger negative responses and non-rumours tend to receive positive ones; cumulative emotion regression coefficient showed that negative emotions grow significantly faster in rumours comments as positive emotions in non-rumour ones; the rumour tweets are not directly connected with other emotions and rely on the emotion surprise as a bridge. Pessimism is primarily influenced by negative emotions (sadness and fear), while optimism is causally influenced by positive emotions (joy, love, and trust), anger and disgust exhibit bidirectional causation. By presenting the framework, we hope to facilitate research in more comprehensive and fine-grained study in emotion in online rumour contents and better detection techniques.
@@
This work also faces several challenges and limitations. (1) We rely on EmoLLM as our automatic emotion detection tool for all emotion-related tasks. While it is generally efficient and effective, it exhibits inaccuracies in analyzing complex online discussions, such as those involving sarcasm. (2) Although we have access to the chronological order of tweets within conversations, explicit conversation structures (i.e. the reply-to structure) are not available for all data. (3) The datasets used in this study are limited to English textual rumour data. Future work should explore multilingual and multimodal content in rumour conversations to provide a more comprehensive analysis. (4) The choice of datasets was also impacted by more restricted access to current social platform APIs and the limited availability of suitable, publicly accessible datasets. However, we will prioritize efforts to replicate our framework on newer datasets as they become accessible.
@@
is under CC-BY license, Twitter15 and Twitter 16666https://github.com/majingCUHK/rumour_RvNN are under MIT License, EmoLLM [^7] https://github.com/lzw108/EmoLLMs is under MIT License and Huggingface Transformers [^8] https://github.com/huggingface/transformers is under Apache License 2.0). Our source code and annotated data will be under MIT license.
Annotation Guideline
Table 6 outlines the guidelines used by annotators for evaluating multi-label emotion classification, sentiment valence, and emotion intensity. Due to the cognitive load of annotating all emotions for intensity, we focus on fear—a common emotion in both rumour and non-rumour threads—for human evaluation. Our annotation team comprised three in-house researchers with varying levels of experience: one PhD candidate with expertise in various NLP tasks and social media text analysis, one Master’s student with a linguistics background, and one Bachelor’s student who was in the early stages of research and new to the area. While not ideal, this setting aims to approximate the range of comprehension and analytical abilities present within a typical online audience encountering rumors on social media.
Annotation Guideline
Multi-label Emotion Annotation: Categorize the text’s emotional tone as either ‘neutral or no emotion’ or identify the presence of one or more of the given emotions (anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust), emotions are separated by comma (,), no period at the end, e.g., anger, fear.
Sentiment Valence Annotation: Categorize the text into an ordinal class that best characterizes the writer’s mental state, considering various degrees of positive and negative sentiment intensity. 3: very positive mental state can be inferred. 2: moderately positive mental state can be inferred. 1: slightly positive mental state can be inferred. 0: neutral or mixed mental state can be inferred. -1: slightly negative mental state can be inferred. -2: moderately negative mental state can be inferred. -3: very negative mental state can be inferred
Emotion Intensity (Fear): Assign a categorical value from 0 to 3 to represent the intensity of emotion fear expressed in the text: – 0 - No such emotion detected – 1 - Minor expression of the emotion – 2 - oderate expression – Strong expression of the emotion
Table 6: Annotation Guideline for Multi-label emotion classification, Sentiment Valence and Emotion Intensity annotation.